An open box is to be made with a capacity of 36 000 cu cm. Find the least material

used in making the box if the length is twice its width.


1
Expert's answer
2021-05-07T10:11:47-0400

Answer:-

"l=length\\width = w\\h=height"


given l = 2w


"Volume(V) = l\times w \times h"

"V = 2w\times w \times h \\
V = 2w^2\times h
\\36000=2w^2\times h\\
now \space for \space the \space material \\

Area = lw + 2wh + 2lh \\
= 2w^2 + {36000\over w} + {72000\over w}
\\
=2w^2+{108000\over w }\\
now\\
by\space calculus \\
\frac{d(area)}{dw} = 4w - {108000\over w^2}\\
now\\
\frac{d(area)}{dw}=0\\
clearly \ w = 30 cm
\\
and \ now \ l=2w \ so \ , \ l=60 cm \ h = 20cm\\

Area = lw + 2wh + 2lh \\
\boxed{Area = 5400 \ sq \ cm}answer"